OpenStudy (anonymous):

(2/3j)to the power of 2

OpenStudy (redheadangel):

just multiply 2/3 by 2/3rds

OpenStudy (redheadangel):

then put the j after the answer

OpenStudy (anonymous):

= (2/3j) x (2/3j) = (2 x 2) / (3j x 3j) Not sure if j is a complex number, if its not then = 4/9j^2 If it is complex then = -4/9

TheSmartOne (thesmartone):

\[(\frac{ 2 }{ 3j })^2=\frac{ 2^2 }{ (3j)^2 }=\frac{ 4 }{9j^2}\]

TheSmartOne (thesmartone):

@sd2452 Are you sure there is a J there?

OpenStudy (redheadangel):

so the answer is: 4/9j

TheSmartOne (thesmartone):

j^2..